OpenStudy (anonymous):

2. Solve the following system. Show all your work. {3x - 2y + 2z = 30 {-x + 3y - 4z = -33 {2x - 4y + 3z = 42

OpenStudy (texaschic101):

do you know the elimination method ?

OpenStudy (anonymous):

know ABOUT it yes. but i don't really understand it all to well.

OpenStudy (texaschic101):

take any two of your equations and try to eliminate a variable. Sometimes you have to multiply an equation by a certain number to make the variable cancel out. 3x - 2y + 2z = 30 -x + 3y - 4z = -33 -->(3) ---------------- 3x - 2y + 2z = 30 -3x + 9y - 12z = - 99 (result of multiplying by 3) ----------------add 7y - 10z = - 69 do you understand what I did so far ?

OpenStudy (anonymous):

yeah!

OpenStudy (texaschic101):

now we will take the equation we didn't use and either of the other equations and eliminate x as well. Whatever variable we eliminate in the first two equations, we have to eliminate in the last two equations. So we have to eliminate x. -x + 3y - 4z = -33 -->(2) 2x - 4y + 3z = 42 ----------------- -2x + 6y - 8z = - 66 (result of multiplying by 2) 2x - 4y + 3z = 42 ----------------add 2y - 5z = - 24 you still with me ? Have any questions so far ?

OpenStudy (anonymous):

yep! none so far!

OpenStudy (texaschic101):

Now we take the " answers " we got (the 2 equations that we eliminated x in), and add them together eliminating a variable. 7y - 10z = -69 2y - 5z = -24 -->(-2) -------------- 7y - 10z = -69 -4y + 10z = 48 (result of multiplying by -2) -------------add 3y = -21 (divide by 3) y = -7 Now we have solved a variable. You understanding so far ?

OpenStudy (anonymous):

WOW! yep! your a great teacher!! I'm understanding it clearer now! Thanks!

OpenStudy (texaschic101):

now we will sub -7 in for y in one of the equation that we eliminated x in. 7y - 10z = -69 7(-7) - 10z = -69 -49 - 10z = -69 -10z = -69 + 49 -10z = - 20 z = 2 catching on now ?

OpenStudy (anonymous):

yep

OpenStudy (texaschic101):

now that we know y and z, we will sub those answers into one of the original equations to find x. -x + 3y - 4z = -33 -x + 3(-7) - 4(2) = -33 -x - 21 - 8 = -33 -x - 29 = -33 -x = -33 + 29 -x = - 4 x = 4 and now we will check our answers by subbing in what we got for the variables into one of the original equations. I suggest checking your answers because in elimination, mistakes can be easily made. Lets check... 2x - 4y + 3z = 42 2(4) - 4(-7) + 3(2) = 42 8 + 28 + 6 = 42 14 + 28 = 42 42 = 42 (correct) if it comes out equal, then your answers are correct. x = 4, y = -7 and z = 2 any questions ?

OpenStudy (anonymous):

nope you answered my question completely! thanks!

OpenStudy (texaschic101):

no problem :) I hope I made it easier for you to understand

OpenStudy (anonymous):

Yes you did! Thanks again for taking the time!

OpenStudy (texaschic101):

thats what I am here for :)